I'm suppose to simplify e raised to ln2^5 and i moved the 5 in front of the ln but i don't know how to simplify it further, is there any more simplifications that can be made
Found 2 solutions by solver91311, DrBeeee:
Answer by solver91311(24713)





and , so...

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by DrBeeee(684)

You have
(1) x = e^(ln2^5)
Take the natural logarithm of each side to get
(2) ln(x) = (ln2^5)*ln(e)
Since ln(e) = 1 we have
(3) ln(x) = ln(2^5)
Now take the inverse log of each to get
(4) x = 2^5 or
(5) x = 32
Answer: e^(ln2^5) = 32